# A Multi-agent LLM-based JUnit Test Generation with Strong Oracles

### v2

@@ 4.3 Evaluation Metrics and Statistical Testing
This metric measures the proportion of code executed by passing tests in the generated test suites, typically reported as line coverage and branch coverage. Line coverage refers to the percentage of lines that are executed by the test suite, while branch coverage measures the percentage of control flow branches (e.g., if/else conditions) that are reached. Code coverage is a widely used metric for assessing the effectiveness of test prefixes, but it does not address test oracles Hemmati (2015).
@@
This metric measures the accuracy of the assertions (oracles) in a generated test suite. Specifically, it is calculated as the number of correct oracles divided by the total number of oracles in the test suite. An oracle is considered correct if it accurately reflects the intended behavior of the program as specified by its natural language description or reference implementation.
@@
To compare CANDOR against baselines, we perform the Wilcoxon Signed Rank test for all RQs, as recommended in Arcuri and Briand (2011), with a significance level of 0.05. The Wilcoxon Signed Rank test is a non-parametric statistical test used to determine whether there is a significant difference between two independent distributions (e.g., method A vs. method B). For RQ2, we also report the effect size to measure the magnitude of difference in oracle correctness between CANDOR and LLM-Empirical. ranges between 0 and 1, representing the probability that method A yields better results than method B.
4.4 Implementation
All experiments are conducted on a Precision 7960 Tower XCTO workstation equipped with an Intel Xeon w9-3495X processor and dual NVIDIA RTX 6000 Ada GPUs. The implementation is written in Python, using the LangChain library 202 (Accessed: 2025c) for LLM integration. We use LLama 3.1 70B as the basic LLM and DeepSeek R1 LLama-distilled 70B as the reasoning LLM. The max_attempt parameter in the Initialization and Test Prefix Generation step is set to be 3. We limit the number of output tokens of DeepSeek to 2000 for efficiency. The number of “Panelist + Interpreter” pipeline is set to be 3. To support reproducibility, we plan to release the code publicly upon paper acceptance.
5 Experiment Results
5.1 RQ1 Results: Test Prefix Quality
Table 1 presents the experimental results for RQ1, comparing CANDOR with LLM-Empirical and EvoSuite in terms of test prefix quality, measured by line coverage (“Line”), branch coverage (“Branch”), and mutation score (“Mutation”). Overall, CANDOR outperforms both baselines across all metrics, with one exception: on the Leetcode-Medium dataset, EvoSuite achieves slightly higher branch coverage than CANDOR by a margin of 0.01.
@@ 5.1 RQ1 Results: Test Prefix Quality
| LLM-Empirical | 0.885 | 0.930 | 0.910 (2223/2443) | 0.877 | 0.902 | 0.804 (491/611) | 0.853 | 0.927 | 0.901 (774/859) |
| EvoSuite | 0.961 | 0.942 | 0.858 (2096/2443) | 0.959 | 0.959 | 0.845 (516/611) | 0.984 | 0.976 | 0.888 (763/859) |
| CANDOR | 0.991* | 0.970 | 0.980* (2384/2443) | 0.990* | 0.949 | 0.939* (574/611) | 0.989* | 0.980 | 0.937* (805/859) |
@@
For line coverage, CANDOR consistently achieves the highest scores across all datasets, with a minimum of 0.989 on the Leetcode-Hard dataset. Wilcoxon Signed Rank tests confirm that the differences between CANDOR and the strongest baseline, EvoSuite, are statistically significant. Similarly, for mutation score, CANDOR also performs best across the board, successfully killing 2384 (out of 2443), 574 (out of 611), and 805 (out of 859) mutants on HumanEvalJava, Leetcode-Medium, and Leetcode-Hard, respectively. Wilcoxon tests also confirms the significance of these improvements over EvoSuite. For branch coverage, although CANDOR achieves the best results on two out of three datasets, none of the differences between CANDOR and EvoSuite are statistically significant.
@@
We also observe that CANDOR’s performance remains stable as the dataset shifts from Leetcode-Medium to Leetcode-Hard, with only slight drops in line coverage () and mutation score (), indicating robustness to increasing problem difficulty. However, since both datasets are still method-level and contain limited control-flow complexity, and further evaluation on project-level datasets is needed.
5.2 RQ2 Results: Test Oracle Quality
@@
Figure 4 compares the correctness of oracles generated by CANDOR and LLM-Empirical. We exclude EvoSuite from this comparison since it can only produce regression oracles rather than specification-based oracles.
CANDOR achieves oracle correctness scores of 0.971, 0.993, and 0.961 on HumanEvalJava, Leetcode-Medium, and Leetcode-Hard, respectively. CANDOR substantially outperforms LLM-Empirical on all datasets, with the smallest margin of 0.158 () on HumanEvalJava and the largest margin of 0.251 () on Leetcode-Hard. Wilcoxon Signed Rank tests confirm all differences are statistically significant, and the effect size () indicates a strong advantage.
To evaluate performance on faulty code, we also test on mutants generated by PiTest. Both methods show a decrease in oracle correctness, but CANDOR ’s decline is smaller, widening the gap with LLM-Empirical. The gaps on HumanEvalJava, Leetcode-Medium and Leetcode-Hard are 0.168 (), 0.184 () and 0.186 (0.952-0.766), respectively. This is likely because CANDOR leverages both source code and natural language descriptions for oracle generation, placing greater emphasis on the latter. As a result, errors in the source code have limited impact on CANDOR, since the description-derived requirements help correct the oracle. Wilcoxon tests again confirm significance of all the differences, with an even stronger effect size ().
5.3 RQ3 Results: Ablation study
@@
| CANDOR | 0.991* | 0.970* | 0.980* (2384/2443) | 0.971* | 0.990* | 0.949* | 0.939* (574/611) | 0.993 | 0.989* | 0.980* | 0.937* (805/859) | 0.961* |
| w/o Planner | 0.892 | 0.840 | 0.869 (2125/2443) | N/A | 0.940 | 0.903 | 0.869 (531/611) | N/A | 0.935 | 0.922 | 0.833 (716/859) | N/A |
| w/o Req. | N/A | N/A | N/A | 0.939 | N/A | N/A | N/A | 0.988 | N/A | N/A | N/A | 0.920 |
| w/o Panel | N/A | N/A | N/A | 0.873 | N/A | N/A | N/A | 0.941 | N/A | N/A | N/A | 0.901 |
@@
Removing the Planner significantly reduces test prefix quality across all datasets. The minimum decreases are observed on Leetcode-Medium with line coverage, branch coverage, and mutation score dropping by at least 0.050 (), 0.046 (), and 0.070 (), respectively. Wilcoxon tests confirm these decreases are statistically significant. This underscores the critical role of the Planner, which analyzes current coverage and strategically plans to cover missing lines.
Removing the Requirement Engineer or the panel discussion primarily impacts oracle correctness. Without the Requirement Engineer, oracle correctness decreases slightly: 0.022 (), 0.005 () and 0.041 () on HumanEvalJava, Leetcode-Medium, and Leetcode-Hard, respectively. Wilcoxon tests shows these changes on HumanEvalJava and Leetcode-Hard are significant while the decline on Leetcode-Medium is not. This likely suggests that the descriptions of source code in these three datasets are clear and well-structured, which reduces the need for explicit requirement parsing using the Requirement Engineer agent. However, in real-world scenarios with less clear documentation, this agent is expected to provide greater benefits.
In contrast, removing the panel discussion causes substantial and statistically significant drops in oracle correctness—0.098 (), 0.052 (), and 0.060 () across datasets. This aligns with expectations, as without the panel discussion, LLMs are more prone to uncertainty and hallucination, generating plausible but incorrect oracles more frequently.
6 Discussion
Despite the high effectiveness of incorporating LLMs for unit test generation, instructing LLMs to act as expected is a challenging task. In this section, we share several key insights obtained from our experiments about effective LLM instruction (Section 6.1), including the adoption of specialized agents, the strategy to mitigate LLM hallucination and the strategy to handle the verbosity of reasoning LLMs. We also report the limitations of our method in Section 6.2.
6.1 Key Insights
@@
LLMs are increasingly studied not only due to their effectiveness but also their ease of use, as many tasks can be handled by crafting prompts for one single LLM Guo et al. (2024). However, we found these single LLM-based approaches fall short in complex tasks like unit test generation. Unit test generation involves several sub-tasks, such as planning and generating test cases. When an LLM is asked to do both simultaneously, it often gets confused between roles. For example, we observed that when tasked with both planning and generating Java tests, the LLM sometimes produced syntax like [1] instead of the correct Java syntax Arrays.asList(1). We hypothesize this happens due to the interference of the planning sub-task, which is often expressed in natural language or weak typed language like Python in real-world and LLMs’ pretraining corpora. The LLM blends conventions of planning and Java test case generation, resulting in the syntactically wrong test case. By decomposing complex tasks and assigning each sub-task to a specialized LLM agent, we prevent this interference and improve the effectiveness of CANDOR in unit test generation.
@@
LLMs are notoriously prone to hallucination, generating plausible yet nonfactual content Huang et al. (2025). We found that hallucination is particularly frequent in test oracle generations, where LLMs produce incorrect oracles even when clear instructions are present. For instance, when told to square the integer if its index is a multiple of 3, an LLM may incorrectly leave the number unchanged. To address this, we use multiple LLMs to independently generate oracles and then compare their outputs. Since hallucinations are usually inconsistent, taking a consensus across models helps filter them out. Instead of simple majority voting, we let the LLMs explain their reasoning in a panel discussion, and determine the answer with the most consistent logic.
@@
Reasoning LLMs like DeepSeeek R1 are effective but often overly verbose, producing lengthy outputs that significantly increase generation time. In our experiment, using DeepSeek directly for oracle evaluation sometimes resulted in outputs of over tokens, taking hours to complete a single test file. In fact, we observed that correct oracles were often identified early by DeepSeek, but the model would continue second-guessing itself by thinking “I think the correct oracle should be 147. But, wait, maybe I made a mistake ...”. To prevent this “overthinking”, we truncate DeepSeek’s output at tokens and use a basic LLM to extract the correct oracle from its reasoning. This dual-LLM pipeline preserves the benefit of deep reasoning while keeping outputs concise.
6.2 Threats to validity

### v6

@@ 4.3. Evaluation Metrics and Statistical Testing
This metric measures the proportion of code executed by passing tests in the generated test suites, typically reported as line coverage and branch coverage. Line coverage refers to the percentage of lines that are executed by the test suite, while branch coverage measures the percentage of control flow branches (e.g., if/else conditions) that are reached. Code coverage is a widely used metric for assessing the effectiveness of test prefixes, but it does not address test oracles (Hemmati2015HowCriteria).
@@
This metric measures the accuracy of the assertions (oracles) in a generated test suite. Specifically, it is calculated as the number of correct oracles divided by the total number of oracles in the test suite. An oracle is considered correct if it accurately reflects the intended behavior of the program as specified by its natural language description or reference implementation. Specifically in our experiments, both HumanEvalJava and LeetCodeJava are well-established projects with no known bugs. Therefore, any failure of a generated test oracle suggests it is incorrect.
@@
To compare CANDOR against baselines, we perform the Wilcoxon Signed Rank test for all RQs, as recommended in (Arcuri2011AEngineering), with a significance level of 0.05. The Wilcoxon Signed Rank test is a non-parametric statistical test used to determine whether there is a significant difference between two independent distributions (e.g., method A vs. method B). For RQ2, we also report the effect size to measure the magnitude of difference in oracle correctness between CANDOR and LLM-Empirical. ranges between 0 and 1, representing the probability that method A yields better results than method B.
4.4. Implementation
All experiments are conducted on a Precision 7960 Tower XCTO workstation equipped with an Intel Xeon w9-3495X processor and dual NVIDIA RTX 6000 Ada GPUs. The implementation is written in Python, using the LangChain library (2025LangchainWebsite) for LLM integration.
Most works in the literature use EvoSuite as a baseline with its default, recommended configuration (“assertion_
strategy=‘‘mutation’’; assertion_timeout=60s”) (Tufano2021UnitContext; chen2024chatunitest; Hossain2024TOGLL:LLMs; Siddiq2024UsingStudy; yuan). However, to ensure EvoSuite is fully exercised, we experimented with different time budgets (1, 2, 3, 4, 5, 10, 30, 60, and 120 minutes per SUT) and assertion strategies (”MUTATION”, ”ALL”, and ”UNIT”). We find the resulting coverage was not further improved after “assertion_timeout 2 min”, stuck at a plateau as previously reported (lemieux2023codamosa). The choice of assertion strategies did not significantly affect the quality of the generated tests. Therefore, we utilize EvoSuite with the default assertion strategy (assertion_strategy=‘‘mutation’’) and an increased timeout threshold (assertion_timeout=2 min). For PiTest, we employ the default group of mutation operators following the common practice in the literature (shi2019mitigating; de2020using; wang2025towards).
We use Llama 3.1 70B as the basic LLM and DeepSeek R1 Llama-distilled 70B as the reasoning LLM. DeepSeek R1 is selected for its strong reasoning capability among open-source models. For the basic LLM, we also experimented with other commonly used alternatives such as CodeLlama 70B and Mistral 22B, however, Llama 3.1 70B achieved the best overall performance. The corresponding results are reported in the Appendix for completeness. The max_attempt parameter in the Initialization and Test Prefix Generation step is set to 3. If CANDOR fails to generate an initial test file for a given method under test within these attempts, it stops further trials and proceeds to the next method. We limit the number of output tokens of DeepSeek to 2000 for efficiency. The number of “Panelist + Interpreter” pipeline is set to be 3. We experimented with 1 to 5 pipelines. However, the improvements beyond 3 are not significant and incur much longer generation time. To support reproducibility, we plan to release the code publicly upon paper acceptance.
5. Experiment Results
5.1. RQ1 Results: Test Prefix Quality
Table 1 presents the experimental results for RQ1, comparing CANDOR with LLM-Empirical and EvoSuite in terms of test prefix quality, measured by line coverage (“Line”), branch coverage (“Branch”), and mutation score (“Mutation”).
@@ 5.1. RQ1 Results: Test Prefix Quality
| LLM-Empirical | 0.704* | 0.688* | 0.823* (2011/2443) | 0.771* | 0.732* | 0.868* (530/611) | 0.714* | 0.729* | 0.778* (668/859) |
| EvoSuite | 0.961* | 0.942 | 0.858* (2096/2443) | 0.959* | 0.959 | 0.845* (516/611) | 0.984 | 0.976 | 0.888* (763/859) |
| CANDOR | 0.991 | 0.950 | 0.980 (2384/2443) | 0.990 | 0.949 | 0.939 (574/611) | 0.989 | 0.980 | 0.937 (805/859) |
@@
CANDOR consistently achieves high line coverage across all datasets, with the lowest value still reaching 0.989 on the Leetcode-Hard dataset. Compared to EvoSuite, CANDOR shows marginal improvements, with a maximum gain of 0.031 ( on Leetcode-Medium). In contrast, LLM-Empirical performs substantially worse than CANDOR across all three datasets, with the largest gap reaching 0.287 () on HumanEvalJava. Wilcoxon Signed Rank tests confirm that the differences between CANDOR and LLM-Empirical are statistically significant (), whereas the differences between CANDOR and EvoSuite are not statistically significant ().
For branch coverage, both CANDOR and EvoSuite substantially outperform LLM-Empirical. The maximum branch coverage achieved by LLM-Empirical is 0.732 on Leetcode-Medium, whereas the minimum branch coverage achieved by EvoSuite or CANDOR is 0.942 on HumanEvalJava. Wilcoxon Signed Rank tests confirm that the differences between CANDOR and LLM-Empirical are statistically significant (), whereas the differences between CANDOR and EvoSuite are not statistically significant ().
For mutation score, CANDOR performs the best across all three datasets, successfully killing 2384 (out of 2443), 574 (out of 611), and 805 (out of 859) mutants on HumanEvalJava, Leetcode-Medium, and Leetcode-Hard, respectively. CANDOR outperforms EvoSuite by at least 0.049 ( on Leetcode-Hard) and LLM-Empirical by at least 0.157 ( on HumanEvalJava). Wilcoxon Signed Rank tests confirm that all differences between CANDOR and the baselines are statistically significant (). We hypothesize that the advantage of CANDOR in mutation scores stems from the LLM’s semantic understanding. Unlike EvoSuite, which primarily evolves test prefixes to satisfy structural coverage criteria, CANDOR also considers the semantics of the program under test (e.g., respecting input constraints, maintaining data invariants, and producing realistic usage patterns). These semantically meaningful test prefixes are more effective at revealing subtle behavioral deviations introduced by mutants, even when overall coverage is similar. Consequently, CANDOR achieves substantially higher mutation scores than EvoSuite.
@@
We also observe that CANDOR’s performance remains stable as the dataset shifts from Leetcode-Medium to Leetcode-Hard, with only slight drops in line coverage () and mutation score (), indicating robustness to increasing problem difficulty. We posit that this robustness originates from the pretraining of LLMs on a vast and diverse corpus, which allows CANDOR to generalize well across programs of varying complexity in this context.
Overall, CANDOR outperforms both baselines across all metrics, with one exception: on the Leetcode-Medium dataset, EvoSuite achieves slightly higher branch coverage than CANDOR by a margin of 0.01. However, the differences between CANDOR and EvoSuite in terms of line and branch coverage are marginal and mostly insignificant, indicating that CANDOR achieves comparable performance in code coverage to EvoSuite. Moreover, CANDOR significantly outperforms EvoSuite in mutation scores, suggesting better bug-detection capability.
5.2. RQ2 Results: Test Oracle Quality
@@
Figure 4 compares the correctness of oracles generated by CANDOR, LLM-Empirical, and TOGLL. We exclude EvoSuite from this comparison since it can only produce regression oracles rather than specification-based oracles.
On Correct Code. As depicted in the left part of Figure 4, CANDOR achieves the highest oracle correctness scores of 0.910, 0.930, and 0.894 on HumanEvalJava, Leetcode-Medium, and Leetcode-Hard, respectively. The gap between CANDOR and LLM-Empirical is substantial, with a minimum of 0.467 () on HumanEvalJava in terms of oracle correctness. When compared to TOGLL, we refer to EVO-CANDOR’s results since they both use the same EvoSuite test prefixes. We observe that, despite the slight performance decrease compared to CANDOR, EVO-CANDOR still outperforms TOGLL by a large margin, with a minimum of 0.255 () on Leetcode-Hard. We conduct the Wilcoxon Signed Rank tests for all the comparisons, and the results confirm that all differences are statistically significant (). We further calculate Vargha and Delaney’s effect size for the comparison between EVO-CANDOR and TOGLL and found a strong effect size (), indicating there is a chance that EVO-CANDOR can outperform TOGLL.
We posit that the superior performance of CANDOR can be explained by its capability to generate oracles based on only informal descriptions. In contrast, the optimal configuration of TOGLL requires not only the descriptions but also method signatures, which are often unavailable in real-world projects and are absent in our datasets. Moreover, TOGLL is fine-tuned on the SF110 dataset, consisting of tests generated with EvoSuite v1.0.6, Java 8, and Ant Builder. Program features beyond this environment cannot be recognized, thereby diminishing the effectiveness of TOGLL. For instance, we observe that TOGLL struggles with newer Java features, such as “var” for type inference and lambda expressions such as “list.stream().map(String::toUpperCase).toList();”
On Faulty Code. As described in Section 4, the faulty code is created by PiTest, which introduces mutations into correctly-implemented benchmarks: HumanEvalJava, Leetcode-Medium, and Leetcode-Hard. As shown in the right part of Figure 4, CANDOR remains the best approach across all datasets. Compared to LLM-Empirical, the gaps are even larger than those in correctly-implemented code in terms of oracle correctness, with a minimum of 0.506 () on Leetcode-Medium. When comparing with TOGLL, EVO-CANDOR outperforms it by a large margin, with a minimum of 0.211 () on Leetcode-Medium. Wilcoxon Signed Rank tests confirm that all differences in the comparisons between CANDOR and LLM-Empirical, and between EVO-CANDOR and TOGLL, are statistically significant (). We also observe an even stronger Vargha and Delaney’s effect size () in the comparison between EVO-CANDOR and TOGLL, as there is a chance for EVO-CANDOR to outperform TOGLL.
The superior performance of CANDOR on faulty code highlights its fault-detection capability. We believe such performance is achieved by deriving oracles from natural language descriptions of the source code in an effective manner, rather than relying solely on the source code itself. Specifically, the panel discussion employed by the CANDOR enables the cross-validation on the generated oracles. This design prevents hallucination, yielding more accurate oracles than approaches that rely on single-agent reasoning.
Again, it is important to recall that comparing CANDOR with TOGLL is somewhat unfair, as TOGLL is fine-tuned with an extra dataset, while CANDOR uses off-the-shelf LLMs. Despite this advantage, CANDOR still outperforms TOGLL substantially while offering greater flexibility. Indeed, unlike TOGLL, CANDOR is flexible when applied to new Java versions or system environments, eliminating the need for fine-tuning on a version-specific dataset. This is particularly important in practice as many software systems and environments are frequently upgraded, making fine-tuning costly and time-consuming.
5.3. RQ3 Results: Ablation study
@@
| CANDOR | 0.991 | 0.970 | 0.980 (2384/2443) | 0.910 | 0.990 | 0.949 | 0.939 (574/611) | 0.930 | 0.989 | 0.980 | 0.937 (805/859) | 0.894 |
| w/o Planner | 0.892* | 0.840* | 0.869* (2125/2443) | N/A | 0.940* | 0.903* | 0.869* (531/611) | N/A | 0.935* | 0.922* | 0.833* (716/859) | N/A |
| w/o Req. | N/A | N/A | N/A | 0.882* | N/A | N/A | N/A | 0.923 | N/A | N/A | N/A | 0.873* |
| w/o Panel | N/A | N/A | N/A | 0.824* | N/A | N/A | N/A | 0.855* | N/A | N/A | N/A | 0.827* |
| w/ Voting | N/A | N/A | N/A | 0.896* | N/A | N/A | N/A | 0.910* | N/A | N/A | N/A | 0.877* |
@@
Removing the Planner significantly reduces test prefix quality across all datasets. The minimum decreases are observed on Leetcode-Medium with line coverage, branch coverage, and mutation score dropping by at least 0.050 (), 0.046 (), and 0.070 (), respectively. Wilcoxon Signed Rank tests confirm these decreases are statistically significant (). This underscores the critical role of the Planner, which analyzes current coverage and strategically plans to cover missing lines. In our experiments, we observe that without the Planner, CANDOR generates tests exploring easy code paths in most cases, lacking a coherent strategy to explore untested code paths.
Removing the Requirement Engineer or the panel discussion primarily affects oracle correctness. In the absence of the Requirement Engineer, oracle correctness decreases slightly by 0.028 (), 0.007 (), and 0.021 () on HumanEvalJava, Leetcode-Medium, and Leetcode-Hard, respectively. Wilcoxon Signed Rank tests indicate the changes on HumanEvalJava and Leetcode-Hard are statistically significant (), whereas the decline on Leetcode-Medium is not significant (). This likely suggests that the descriptions of source code in these three datasets are clear and well-structured, which reduces the need for explicit requirement parsing using the Requirement Engineer agent. However, in real-world scenarios with less clear documentation, this agent is expected to provide greater benefits.
In contrast, removing the panel discussion (“w/o Panel”) causes substantial and statistically significant drops in oracle correctness, i.e., 0.086 (), 0.075 (), and 0.067 () on HumanEvalJava, LeetCodeJava-Medium, and LeetCodeJava-Hard, respectively. Replacing the panel discussion with a majority voting mechanism (“w Voting”) also leads to consistent decreases across all datasets in terms of oracle correctness, with a minimum of 0.014 () on the HumanEvalJava dataset. We performed Wilcoxon tests and found that all decreases are statistically significant (). This aligns with our expectations, as without the panel discussion, LLMs are more prone to uncertainty and hallucination, generating plausible but incorrect oracles more frequently. In our experiments, we observed serious hallucinations during the panel discussion; in over of the cases, the Panelists’ discussions contained clear disagreements. For example, a method “max_element()” that computes the maximum of an array was mistakenly used to compute the minimum instead by one or more Panelists (out of three). However, the Curator can correct these errors by cross-checking the reasoning logic, inputs, and outputs of each Panelist rather than relying solely on majority voting. This design ensures that incorrect or inconsistent outputs are identified and corrected, making the panel discussion more robust than simple majority voting in producing accurate oracles.
6. Discussion
Despite the high effectiveness of LLMs for unit test generation, instructing them to act as expected remains challenging. In this section, we share several key insights from our experiments on effective LLM instruction (Section 6.1), including adopting specialized agents, strategies to mitigate LLM hallucination, and methods to handle the verbosity of reasoning LLMs. We also report the limitations of our method in Section 6.2.
6.1. Key Insights
@@
LLMs are increasingly studied not only due to their effectiveness but also their ease of use, as many tasks can be handled by crafting prompts for one single LLM (Guo2024LargeChallenges). However, we found that these single-LLM approaches fall short on complex tasks, such as unit test generation. Unit test generation involves several subtasks, including planning and generating test cases. When an LLM is asked to perform both tasks simultaneously, it often becomes confused about its roles. For example, we observed that when tasked with both planning and generating Java tests, the LLM sometimes produced syntax like [1] instead of the correct Java syntax Arrays.asList(1). We hypothesize that this happens due to interference from the planning sub-task, which is often expressed in natural language or in weakly typed languages like Python in LLMs’ pretraining corpora. The LLM blends conventions of planning and Java test case generation, resulting in a syntactically wrong test case. By decomposing complex tasks and assigning each sub-task to a specialized LLM agent, we prevent this interference and improve the effectiveness of CANDOR in unit test generation.
@@
LLMs are notoriously prone to hallucination, generating plausible yet nonfactual content (Huang2023AQuestions). We found that hallucination is particularly frequent in test-oracle generations, where LLMs produce incorrect oracles even when clear instructions are provided. For instance, when told to square the integer if its index is a multiple of 3, an LLM may incorrectly leave the number unchanged. To address this, we use multiple LLMs to independently generate oracles and then compare their outputs. Since hallucinations are usually inconsistent, taking a consensus across models helps filter them out. Instead of simple majority voting, we allow the LLMs to explain their reasoning in a panel discussion and determine the answer based on the most consistent logic.
@@
Reasoning LLMs like DeepSeeek R1 are effective but often overly verbose, producing lengthy outputs that significantly increase generation time. In our experiment, using DeepSeek directly for oracle evaluation sometimes produced outputs exceeding tokens, taking hours to complete a single test file. In fact, we observed that DeepSeek often identified correct oracles early. Still, the model would continue second-guessing itself by thinking “I think the correct oracle should be 147. But, wait, maybe I made a mistake ...”. To prevent this “overthinking”, we truncate DeepSeek’s output at tokens and use a basic LLM to extract the correct oracle from its reasoning. This dual-LLM pipeline preserves the benefit of deep reasoning while keeping outputs concise.
6.2. Threats to validity